OpenStudy (anonymous):

In 3 years Dave will be 3 times as old as his brother Ray. A year ago Dave was 7 times as old as Ray. How old are the boys now? So what i have right now for the equations are: D + 3 = 3R 7R - 1 = D Am i correct or at least close? Im usually not very good at these :\ and can i solve it with elimination/substitution or do i use something else?

jimthompson5910 (jim_thompson5910):

Let D = Dave's age now and R = Ray's age now

jimthompson5910 (jim_thompson5910):

In 3 years, Dave will be D+3 years old Ray will be R+3 years old

jimthompson5910 (jim_thompson5910):

So "In 3 years Dave will be 3 times as old as his brother Ray" means that D+3 = 3*(R+3)

jimthompson5910 (jim_thompson5910):

A year ago, Dave was D-1 years old. Ray was R-1 years old So if "A year ago Dave was 7 times as old as Ray", then D-1 = 7*(R-1)

OpenStudy (anonymous):

okay, following so far, so i cant just use elimination on this then?

jimthompson5910 (jim_thompson5910):

You can if you want, but substitution might be easier

jimthompson5910 (jim_thompson5910):

D+3 = 3*(R+3) D+3 = 3R+9 D = 3R+9-3 D = 3R+6

jimthompson5910 (jim_thompson5910):

Now plug this into the second equation D-1 = 7*(R-1) 3R+6-1 = 7*(R-1) and solve for R

OpenStudy (anonymous):

okay, thank you, im going to try to work it out and see if i get it right!

jimthompson5910 (jim_thompson5910):

alright, tell me what you get

OpenStudy (anonymous):

Ray is 3 so far...

OpenStudy (anonymous):

and Dave is 15?

jimthompson5910 (jim_thompson5910):

you are 100% correct

OpenStudy (anonymous):

Yay! Thank you so much! I think i get it now :')

jimthompson5910 (jim_thompson5910):

Check: 3 years from now, Dave is 15+3 = 18 3 years from now, Ray will be 3+3 = 6 Notice how Dave is 3 times as old as Ray ----------------------------- 1 year ago, Dave is 15-1 = 14 1 year ago, Ray was 3-1 = 2 and notice how Dave is 7 times as old as Ray

jimthompson5910 (jim_thompson5910):

So the answer is verified

OpenStudy (anonymous):

Thank you! You really helped me out a lot! i was really stressed over this!

jimthompson5910 (jim_thompson5910):

you're welcome, I'm glad it's making sense now